Wellington Cabinets has fixed costs totaling $96,000. Its contribution margin per unit is $1.50, and the selling

price is $5.50 per unit. If they desire a profit of $150,000, how many units must they sell? How many sales

dollars must they earn?
Answer #1

Required unit = (Fixed cost+Desired profit)/Contribution margin per unit = (96000+150000)/1.5 = 164000 Units

Required sales = 164000*5.5 = $902000